OpenStudy (anonymous):

find the slope and y intercept of the line whose equation is -6x+y=-4

OpenStudy (anonymous):

Let us make this y = mx + b form, to do this we need to get y by itself on 1 side of the equation. We can start by adding 6x, that leaves us with y = 6x - 4. In the structure y = mx + b, m is the slope and b is the y intercept. Thus in y = 6x - 4, the slope is 6 and the y intercept is (0,-4) as the coordinate pair system works as: (x, y)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in y=mx+b m is the slope and b is the y intercept lets turn -6x+y=-4 into y=mx+b by having y alone on one sides:\[-6x+y=-4\]add -6x:\[y=6x-4\]so the slope (m) is 6 and the y intercept (b) is -4

OpenStudy (anonymous):

y=6x-4 y=mx+c slop m=6